You need to construct a 100-pF capacitor for a science project. You plan to cut two L x L metal squares and place spacers between them. The thinnest spacers you have are 0.2 mm thick. What is the proper value of L? If this capacitor is connected to a 9 V battery and charged up, how much charge is on each plate?

The answers are 4.8 cm, 90 nC. I got the first answer, though the second answer I am two decimal places off and cannot figure out why. Thank you
Answer #1

a)

Capacitance of the capacitor is

C=eoA/d

=>A=(100*10-12)(0.2*10-3)/(8.85*10-12)

A=2.26*10-3 m2

Since A=L2=2.26*10-3

L=0.0475 m=4.75 cm

b)

Charge

Q=CV =(100*10-12)(9)

Q=0.9*10-9C =0.9 nC